OpenStudy (anonymous):

Whats the formula for doing slope with only this. (5,0) and (6,8)?

OpenStudy (anonymous):

\[ m = \frac {y_2 - y_1} {x_2 - x_1} \] The above is the slope through the line passing through two points (x₁,y₁) and (x₂,y₂).

OpenStudy (anonymous):

okay... thnk you.. :)\

OpenStudy (amistre64):

yep; subtract the points and stack y over x x , y ( 5 , 0 ) -( 6 , 8 ) --------- -1, -8; slope = y/x = -8/-1 = 8

OpenStudy (anonymous):

why are yoiu saying all this stuff is negative?

OpenStudy (amistre64):

it doesnt matter how you subtract the points; in the end it all comes out to a slope of 8 for example: ( 6 , 8 ) -( 5 , 0 ) --------- 1 , 8 ; slope = y/x = 8/1 = 8

OpenStudy (anonymous):

what about this .... (4,-3) and (6,-4)

OpenStudy (amistre64):

what do you get when you subtract one point from the other?

OpenStudy (anonymous):

i am stuck on -4--3

OpenStudy (amistre64):

(4 , 3) -(6 , -4) --------- -2 , 7 right? 4-6 = -2; 3--4 = 7

OpenStudy (anonymous):

but its -4 and -3

OpenStudy (amistre64):

your right, I missed the - :)

OpenStudy (anonymous):

-4 - (-3) = -4 + 3 = -1

OpenStudy (amistre64):

(4 , -3) -(6 , -4) --------- -2 , 1 then :)

OpenStudy (amistre64):

(6 , -4) -(4 , -3) --------- 2 , -1 is just as good

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

i like to stack the points becasue I tend to switch out the x and y values if I do it the other way

OpenStudy (anonymous):

okay i am in an advanced math class and i just get stuck on some

OpenStudy (amistre64):

remember; when ever you get the same sign twice, its a + 4--3 means 4+3 5++2 mean 5+2 whevenever you get opposite signs together, you go negative 6-+2 means 6-2 3+-5 mean 3-5